OpenStudy (broskishelleh):

1. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. using complete sentences, explain whether f (h(x)) and h (f(x)) will always result in the same number. You will use the function f(x) that you created in problem number My Function was: f(x) = (x-1)/4

OpenStudy (broskishelleh):

@Australopithecus

OpenStudy (broskishelleh):

@Compassionate @CoolRanchDoritos

OpenStudy (broskishelleh):

@ikram002p

OpenStudy (broskishelleh):

@OOOPS

OpenStudy (broskishelleh):

@Elsa213 @e.mccormick @Abhisar @am!rah @adilalvi

Elsa213 (elsa213):

yes?

OpenStudy (broskishelleh):

May I have assistance?

Elsa213 (elsa213):

O ok

Elsa213 (elsa213):

So first of all, what you just gave me as a function, isn't really a function. It's an equation with the solution x=4. So as a function, let's use f(x)=4x+2. To plug in 3, all you have to do, is replace the x with 3. So \(f(3)=f(3)+2=12+2=14\).

Elsa213 (elsa213):

So now we create a new function \(h(x)\). For the sake of argument, let's suppose this function is \(2x+6\). Using this new function, can you tell me what \(h(3)\) is?

Elsa213 (elsa213):

So \(f(h(3))=f(12)\). Plugging 12 into \(f(x)\), we get \(f(12)=4(12)+2=48+2=50\). Repeating the same process, can you tell me what \(h(f(3)\) is? (recall that \(f(3)=14)\)

Elsa213 (elsa213):

Well just by plugging in 14 to h(x), we get h(14)=2(14)+6=28+6=34. Now when we compare f(h(3)) and h(f(3)), we notice that f(h(3))=50 and h(f(3))=34. These numbers are not equal. So this means that f(h(x)) and h(f(x)) are NOT always the same number.

OpenStudy (broskishelleh):

12 + 6 = 18

OpenStudy (broskishelleh):

f(3) = 18

Elsa213 (elsa213):

yerp

Elsa213 (elsa213):

@broskishelleh did u understand?